Monica has decided that she wants to build enough retirement wealth that, if invested at 7 percent per year, will provide her with $3,000 monthly income for 30 years. To date, she has saved nothing, but she still has 20 years until she retires. How much money does she need to contribute per month to reach her goal?
Answer #1

Step 2: FV = 450922.70, I = 7/12 = 0.5833, N = 20 × 12 = 240, PV = 0, CPT PMT = −865.62.



answered by: Kraer4342